Suppose a security with a risk-free cash flow of $152 in one year trades for $137 today. If there are no arbitrage opportunities, what is the current risk-free interest rate?
Answer #1

step 1) 137= 152/ (1+r)


step2) 152/ 137= 1+r

1+r= 1.1094890511

1- 1.1094890511= r

r= 0.1094890511

Answer: r= 10.95


answered by: Andrew San Andres